OpenStudy (anonymous):

find the slope of the line: 4y-5=6x

OpenStudy (anonymous):

All we have to do is put the equation into the slope-intercept form: y = mx + b 4y = 6x + 5 -> y = (3/2)x + 5/4 So, your slope is "m"

OpenStudy (anonymous):

Just match up that last equation with the form y = mx + b. Hint: It's the number (fraction) in front of the "x".

OpenStudy (anonymous):

All good now?

OpenStudy (anonymous):

gotcha. haha it took me a minute to get the hint but omg you're the best!

OpenStudy (anonymous):

I'm just on a good streak. You're just good to work with!

OpenStudy (anonymous):

And thx again. I hope you are enjoying your math, or at least as much as one can! lol

OpenStudy (anonymous):

no problem!